hi can you please help me
2
9x4 ---=
3

Answer by Alan3354(69443)

Is it 9 times 4 2/3?
If so,
9*4 = 36
9*(2/3) = 6
36 + 6 = 42
------------
or 9 times 4 2/3 = 9*(14/3)
= 126/3
= 42